OpenStudy (anonymous):

[9.05] Part 1: Create your own quadratic equation that cannot be solved by factoring, but can be solved using the quadratic formula. Identify the values of a, b, and c, and find the solutions using the quadratic formula. Show all work to receive credit. Part 2: Using complete sentences, explain how you know that the equation from Part 1 cannot be solved by factoring, but can be solved by using the quadratic formula.

OpenStudy (anonymous):

x^2 + 3x + 1 is one example. Although there are many A = 1 B = 3 C = 1

OpenStudy (anonymous):

can you finish it from there?

OpenStudy (anonymous):

not really i have till midnight to finish this and i have no idea how to do these

OpenStudy (anonymous):

do you know the quad formula? i would prefer not to type it out

OpenStudy (anonymous):

so write down that formula. Then take your A B and C i gave you. Plug into that formula for a b and c

OpenStudy (anonymous):

no i dont it just want simple instructions

OpenStudy (anonymous):

\[-3 \pm (\sqrt{9 - 4(1)(1)})/2\]

OpenStudy (anonymous):

\[\frac{ -b \pm \sqrt{b ^{2}-4ac} }{ 2a}\] this is the formula a: is x^2 coefficient b: is x coefficient c: is the number